Evaluate the following expressions when: x = 3, y = 5, z = -6
3x+y

Answer by praseenakos@yahoo.com(507)

Question:

Evaluate the following expressions when: x = 3, y = 5, z = -6
3x+y

Answer:

Substitute the value, x = 3, y = 5, z = -6 in the given expression....

==> 3x + y = 3 * 3 + 5 = 9 + 5 = 14.
There is no z term in the given expression, so no need to substitute the value of z.

Hope you understood....

Regards.

Praseena.